OpenStudy (maddog3354):

Perform the following food service calculation. Meat to be served = fish fillet Amount of stock = 86 lb. How many customers can be served fish?

OpenStudy (mathmale):

Have you thought about this situation? Have you considered what portion size is sufficient for one customer? Without this information, without your assuming numerical value(s) for portion size, this question, as stated, really is unanswerable.

OpenStudy (marieblossom16):

you need more information to solve this problem.

OpenStudy (crabbyoldgamer):

If each person eats 86 lbs of fish, you can serve 86/86 = 1 customer. If each person eats 43 lbs of fish, i.e. a portion size is 43 lbs, you can serve 86/43 = 2 customers. If each person eats 3 lbs of fish, you can serve 86/3 = 28.667 people: 28 people with a little bit of fish left over. Now you figure out how many people you can serve if each person eats n lbs of fish.